a man is comming down an incline of angle 30.when he walks with speed 2root3.he has to keep his umbrella vertical to protect himself.the speed of rain is 5m/s .At what angle with vertical should he keep his umbrella when he is at rest so not to get wet

Grade:12

3 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi ,

If you draw the vector diagram of the velocities when the man is in motion, the resultant direcion of the rain is vertical.
Horizontal component of the velocity of man should cancel with that of the rain to make the resultant vertical.

2\sqrt{3}* cos30 = 5 cosa ( a =angle of rain from horizontal)
cosa = 3/5, a = 67 degress

When man is at rest, he should hold his umbrella in the actual direction of rain, which is 67 degress from horizontal.

thanks

yash
9 years ago
th answer should be 37 deegress boss
Shobhit Varshney IIT Roorkee
9 years ago
Hi,
yeah it’s a typing error. The answer should be 53 from horizontal or 37 from vertical.
Do you understand the concept ?

thanks.................................................................................................................